Assume that Home and Foreign produce 2 goods, High-Definition televisions and kiwi fruit. Use the following information to answer the questions.
a) The Home Country is endowed with 10 workers and has the following marginal products: MPLTV = 2, MPLK = 4. Illustrate the Home country’s PPF, placing televisions on the x-axis. Fully label your graph. (2 pts)







b) The Foreign Country is also endowed with 10 workers and has the following marginal products: MPL*TV =1, MPL*K =3. Illustrate the Foreign country’s PPF, placing televisions on the x-axis. Fully label your graph. (2 pts)
  

c) What is the no-trade relative price of TVs in each country? (2 pts)



  
d) Predict the pattern of trade if the world relative price for televisions is PTV/PK = 2.5. (2 pts)

e) What has happened to the real wage for Home country workers? (Be specific) (2 pts)
Answer #1

1).

Consider the given problem here there are two goods “TV” and “fruit”. Now, the marginal productivity of two goods are given by “MPLTV=2” and “MPLK=4” respectively, => there labor requirements are given by, “1/2” and “1/4” respectively. So, the PPF of “Home country” is given by.

=> 0.5*T + 0.25*K = 10, => if we measure “TV” on the “X-axis” and “Fruits” on the “Y-axis”, => the horizontal and the vertical intercepts are given by “10/0.5=20 units” and “10/0.25 = 40 units”. So, the PPF is given by.

Fruit 40 PPF of Home Country 0.5*T 0.25*K 10 20 TV

2).

Now, the marginal productivity of two goods in “Foreign country” are given by “MPLTV* = 1” and “MPLK* = 3” respectively, => there labor requirements are given by, “1” and “1/3” respectively. So, the PPF of “Home country” is given by.

=> 1*T + (1/3)*K = 10, => if we measure “TV” on the “X-axis” and “Fruits” on the “Y-axis”, => the horizontal and the vertical intercepts are given by “10/1 = 10 units” and “10/(1/3) = 30 units”. So, the PPF is given by.

Fruit 30 PPF of Foreign Country T+K/3 10 10 TV

3).

Now, as we know that in autarkic situation the relative price of “TV” is exactly the opportunity cost producing TV. So, for the country the autarkic relative price is “0.5/0.25 = 2” for the home country. Similarly, the opportunity cost of producing TV for the foreign country is given by “1/(1/3) = 3”.

d).

So, here the free trade equilibrium price is “2 < 2.5 < 3”, => home country has lower relative price of TV and foreign country have higher relative price of TV compare to the free trade price, => here under the free trade “home” will export “TV” in exchange of “fruits” and “Foreign country” will export “fruits” in exchange of “TV”.

e)

In home country all the workers will specialize towards the production of “TV”, => the real wage in home country increases.